# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
The electronic ground state of a three-dimensional (3D) band insulator with time-reversal ($\Theta$) symmetry or time-reversal times a discrete translation ($\Theta T_{1/2}$) symmetry is classified by a $\mathbb{Z}_{2}$-valued topological invariant and characterized by quantized magnetoelectric response. Here we demonstrate by explicit calculation in model $\mathbb{Z}_{2}$ topological insulator thin-films that whereas the magnetoelectric response is localized at the surface in the $\Theta$ symmetry (non-magnetic) case, it is non-universally partitioned between surface and interior contributions in the $\Theta T_{1/2}$ (anti-ferromagnetic) case, while remaining quantized. Within our model the magnetic field induced polarization arises entirely from an anomalous ${\cal N}=0$ Landau level subspace within which the projected Hamiltonian is a generalized Su-Schrieffer-Heeger model whose topological properties are consistent with those of the starting 3D model. We identify a novel connection between the ground state geometry of that 3D model and surface-interior-partitioning in thin films.
Introduction— The realization that material properties can depend qualitatively [1, 2] on topological data encoded in the crystal-momentum-dependence of its occupied electronic eigenspace projectors has been an important achievement in modern condensed matter physics. The physical import of this mathematical structure was first recognized in theoretical studies [3, 4] of the integer quantum Hall effect (IQHE) in which the Hall conductivity of two-dimensional band insulators was identified as the (quantized) Chern number of the vector bundle of occupied (electronic) Bloch states over the Brillouin zone (BZ) torus [5]. Subsequently other, more subtle, topological invariants of band insulating ground states have been found to be physically relevant. Among these, three-dimensional (3D) band insulators with time-reversal symmetry (TRS, denoted as $\Theta$) or generalized TRS ($\Theta T_{1/2}$), which have a quantized magnetoelectric response [6, 7, 8],
@@
have been a prominent focus. In Eq. (1) superscript indices identify Cartesian components, ${\mathbf{P}}$ and ${\mathbf{M}}$ are the (electronic) polarization and orbital magnetization, and ${\mathbf{E}}$ and ${\mathbf{B}}$ are uniform dc electric and magnetic fields. In both non-magnetic ($\Theta$ symmetric) and anti-ferromagnetic (AFM) ($\Theta T_{1/2}$ symmetric [9]) cases, $\alpha=0$ or $e^{2}/2hc$ (mod $e^{2}/hc$) and is determined, in part, by the $\mathbb{Z}_{2}$ classification of the ground state [6, 7, 8]. Indeed, an explicit expression has been proposed [6] that relates $\alpha$ to the bulk BZ integral of the Chern-Simons 3-form in the space of occupied Bloch states. The value of $\alpha$ implied by this bulk expression ($\alpha_{\text{CS}}$), quoted explicitly below, is determined modulo $e^{2}/hc$ by the $\mathbb{Z}_{2}$ index of the vector bundle of occupied Bloch states [6, 9, 10, 11].
@@
The relationship between $\alpha_{\text{CS}}$ and the physical magnetoelectric response ($\alpha_{\text{me}}$) of finite volume $\mathbb{Z}_{2}$ topological insulators (TIs) with surfaces is non-trivial [8, 12, 13, 14, 15]. In non-magnetic thin films with surface magnetic dopants that have nonzero surface-normal magnetization components that is opposite on top and bottom surfaces, $\alpha_{\text{me}}$ approaches $\alpha_{\text{CS}}$ in the thick film limit. Here we demonstrate by explicit calculation that whereas the quantized magnetoelectric response is localized at the surface in the non-magnetic case, it is non-universally partitioned between surface and interior contributions in the AFM case. In both cases the response of the (electronic) charge density to magnetic field is local in the sense that it occurs only in the spatial vicinity of static magnetic moments that break $\Theta$ symmetry. The difference is that these are present throughout the film interior in the AFM case. The dipole moment of the charge density that is induced by the magnetic field in the interior region of an AFM film is non-universal, but for sufficiently thick films we find that the sum of interior and surface contributions to the dipole moment is quantized (mod $e^{2}/hc$) at either $\alpha_{\text{me}}=0$ or $e^{2}/2hc$ depending on the $\mathbb{Z}_{2}$ classification of the bulk ground state; in AFM (non-magnetic) thin films (with magnetic surface dopants) $\Theta T_{1/2}$ ($\Theta$) symmetry is broken and therefore, e.g., $\alpha_{\text{me}}=0$ cannot be protected by that symmetry, instead it is protected by the bulk $\Theta T_{1/2}$ ($\Theta$)-symmetry-induced topological classification of the ground state. This property, that the sum of 3D interior and surface contributions is in accord with 3D bulk topological constraints even when the separate contributions are non-quantized and non-universal, is reminiscent of the frequently discussed non-universal Hall current partition between interior and edge [16, 17, 18, 19] in the IQHE.
Our explicit calculations employ a model that can describe both non-magnetic and AFM layered materials, and for each case can support ground states characterized by a trivial or non-trivial $\mathbb{Z}_{2}$ index. The model provides a realistic description of A-type AFM TIs, similar to the paradigmatic AFM TI MnBi2Te4 [20, 21, 22, 23, 24, 25, 26, 27, 28, 29], but is simplified in a way that is convenient for explicit inclusion of magnetic fields that are parallel to the stacking axis [30]. We use a continuum version of this model to calculate electronic charge distributions across thin films and a lattice regularized version to address the topological magnetoelectric response of bulk crystals. As illustrated in Fig. 1 (a), our thin film calculations show that the magnetic field induced dipole moment arises entirely from the model’s anomalous $\mathcal{N}=0$ Landau level (LL) subspace, the dynamics of which are governed by a generalized Su-Schrieffer-Heeger (SSH) Hamiltonian. Because the $\mathcal{N}\neq 0$ LL eigenfunctions are symmetric across the thin film, as illustrated in Fig. 1 (b) and (c), they do not contribute to the magnetic field induced dipole moment. The absence of a dipole moment in the interior of a non-magnetic film is illustrated schematically in Fig. 1 (b), and its presence in an AFM film in Fig. 1 (c).
Anomalous Landau levels—An effective low-energy Hamiltonian for the electronic states in the layered materials of interest can be constructed [31, 32, 30] using Dirac cones, one associated with top and bottom surface of each layer, that are coupled by hybridizing states both within and between layers:
@@
Here we have chosen $\hat{\bm{z}}$ as the stacking axis, ${\mathbf{k}}_{\perp}=(k_{x},k_{y})$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ labels states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$, $\mu,\gamma\in\{\uparrow,\downarrow\}$ are spin labels, $\Delta_{ij}$ are the intra- and inter-layer hybridization parameters, $v_{{}_{D}}$ is the velocity of an isolated Dirac cone, and $m_{i}$ are mass terms that account for exchange coupling to static local magnetic moments (which we take to be aligned with the stacking direction). We will consider only even layer number $N$ so that AFM thin films do not have uncompensated moments, and in our model $m_{i}$ is a mass term at surface $i$ that accounts for exchange coupling to static local magnetic moments. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is non-magnetic and ±$\pm 1$ if layer $l$ is magnetic (for antiferromagnetic configuration we take $M_{l}=(-1)^{l}$); the Dirac-like states at the top or bottom surface of layer $l$ are coupled to the static magnetization of that layer (via $J_{\text{S}}$) and to that of the adjacent layer nearest that surface (via $J_{\text{D}}$), and hybridized only to adjacent surfaces with hopping $\Delta_{\text{S}}$ for opposite surfaces of the same layer and $\Delta_{\text{D}}$ for the closest surface of the adjacent layer (see Fig. 1 of Ref. [30]). In thin film models the Dirac-like states of the outermost surfaces do not have $J_{\text{D}}$ or $\Delta_{\text{D}}$ coupling.
We account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ though the usual prescription  [^1] See, e.g., Chapter 2 of Winkler [56]. of mapping a ${\mathbf{k}}\cdot{\mathbf{p}}$ model to an envelope function approximated (real-space) Hamiltonian that is then minimally-coupled to the electromagnetic field: $\hbar{\mathbf{k}}\rightarrow-i\hbar\bm{\nabla}-(eBx/c)\hat{{\mathbf{y}}}$ in the Landau gauge. Here $e>0$ is the elementary charge unit. Applying this recipe to Eq. (S1) and seeking energy eigenfunctions of the form $\Psi_{E,q_{y}}(x,y)=e^{iq_{y}y}\Phi_{E,q_{y}}(x)/\sqrt{L_{y}}$, where $\Phi_{E,q_{y}}(x)$ has $2N\times 2$ components that correspond with layer and spin label degrees of freedom, results in the individual Dirac Hamiltonian $H_{D}$ of each surface being recast [34] in terms of the LL raising and lowering (differential) operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$,
$H_{D}=\hbar v_{{}_{D}}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}\to\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix}$ (3)
@@
Below we calculate the electric dipole moment of finite thickness films by endowing the LL eigenfunctions of the quasi-2D model Eq. (S1) with $z$-dependence by taking $C^{{\cal N},r}_{(i,\mu)}\rightarrow C^{{\cal N},r}_{(i,\mu)}\delta(z-z_{i})$, where $z_{i}\hat{{\mathbf{z}}}$ is the location of the $i$th surface with the center of the film taken as the origin. The macroscopic electronic polarization (dipole moment per volume) is:
@@
where $\Omega=A(z_{2N-1}-z_{0})$ is the volume of the sample, $A$ is the film cross sectional area and we have noted that only negative energy states are occupied at neutrality. Here the $\Psi_{E}({\mathbf{r}})$ are real-space energy eigenfunctions that have support only within the sample. Taking account of the LL degeneracy $A/(2\pi l_{B}^{2})$ we obtain [34]
@@
The factor in round brackets is the areal electronic charge density on surface $i$ and $d$ is the thickness of the unit cell for each layer. Since $m_{i}$ is odd under the mirror transformation $z\to-z$ [34], it yields a spin-polarization that is odd in $z$. In the ${\cal N}=0$ anomalous LL subspace, which contains only spin-up states, the mass terms give rise to a charge polarization. Because the block projected Hamiltonian is odd under the combined mirror and spin-reversal operations, the charge density contributed by the ${\cal N}\neq 0$ blocks is even under $z\to-z$ and there is no contribution to the dipole moment; the magnetoelectric response of our model comes entirely from the ${\cal N}=0$ block. Since the Hamiltonian of the ${\cal N}=0$ sector is a generalized SSH model with alternating hopping strengths $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$, the quantized magnetoelectric coefficient $\alpha_{\text{me}}$ in our model follows from familiar analogous polarization properties of the SSH model. In the following, we detail how these properties depend on model parameters.
@@
Bulk phase diagram and polarization properties— We can obtain the topological phase diagram of the bulk 3D Hamiltonian related to Eq. (S1) by tracking the band gap evolution when periodic boundary conditions are applied in $\hat{{\mathbf{z}}}$ direction  [^4] Technically, the model presented in Eq. (S1) is a quasi-2D low-energy effective ${\mathbf{k}}\cdot\bm{p}$ Hamiltonian. To generate a genuine 3D model for the electronic states in bulk, we need to replace the ${\mathbf{k}}\cdot\bm{p}$ basis states that the quasi-2D Hamiltonian is expressed in terms of with hybrid Wannier functions that are spatially localized in the ${\mathbf{z}}$ direction and the set of which contains one spin-up and one spin-down WF localized near each surface. We then apply periodic boundary conditions, by which we mean we take the model to describe an infinite chain of coupled-Dirac cone models (i.e., it involves a sum over all lattice vectors) in the ${\mathbf{z}}$-direction.. Our findings are shown in Fig. 2 (a). As explained below, these assignments have been confirmed by calculating the Zak phase of a corresponding one-dimensional (1D) model and also by explicitly evaluating the topological magnetoelectric coefficient of a 3D lattice regularization of Eq. (S1). The model parameters are chosen such that $\delta\equiv J_{\text{D}}/J_{\text{S}}\rightarrow 1$ corresponds to the non-magnetic case, for which a topological phase transition occurs when $|\Delta_{\text{D}}|=\Delta_{\text{S}}$. In this case there is no magnetic-field dependent charge transfer between layers in the bulk as illustrated by the grey curve in Fig. 2 (b). When $\delta\neq 1$ the interior layers are magnetized and magnetic-field dependent charge transfer does occur as illustrated by the green curve in Fig. 2 (b). Note that in the AFM case the topological phase boundary depends on $J_{\text{S}}$, that inter-layer charge transfer occurs in both the trivial ($\mathbb{Z}_{2}$-even) and non-trivial ($\mathbb{Z}_{2}$-odd) phases with magnitude dependent on model parameters, and that the rate of charge transfer with field jumps at the phase transition points, as shown in Fig. 2 (c) and (d).
@@
Figure 2: Topological phase diagram and bulk charge redistribution. (a) Phase diagram of bulk AFM topological insulators vs. $\Delta_{\text{D}}/\Delta_{\text{S}}$ and $\delta\equiv J_{\text{D}}/J_{\text{S}}$ at $\Delta_{\text{S}}/J_{\text{S}}=1$. (The phase boundaries for $J_{\text{S}}=0.8\Delta_{\text{S}}$ and $J_{\text{S}}=1.5\Delta_{\text{S}}$ are plotted as purple and yellow dashed lines.) (b) Schematic magnetic-field induced charge density redistribution for non-magnetic (grey curve) and AFM (green curve) TIs; (c) and (d) Wavefunction amplitude variation (defined in (b)) vs. $\Delta_{\text{D}}$ for various values of $\delta$ when $J_{\text{S}}=1.5\Delta_{\text{S}}$ and $J_{\text{S}}=0.8\Delta_{\text{S}}$, respectively.
@@
The magnetic-field dependent charge transfer seems to suggest that the magnetoelectric responses of both $\mathbb{Z}_{2}$-even and $\mathbb{Z}_{2}$-odd phases are non-zero, but this is not the case. As we have described near Eq. (3) and detailed in the Supplementary Material, the low-energy Hamiltonian in the 3D bulk can be block-diagonalized into subspaces that are uniquely identified by a LL index $\mathcal{N}$. As in finite samples, in this bulk model the magnetoelectric response comes entirely from the $\mathcal{N}=0$ LL subspace, within which the AFM Hamiltonian reduces to a family of identical 1D generalized SSH models – one for each state in the LL (see Supplementary Material):

### v4

# Bulk versus surface: Nonuniversal partitioning of the topological magnetoelectric effect

@@ Abstract
The electronic ground state of a three-dimensional (3D) band insulator with time-reversal ($\Theta$) symmetry or time-reversal times a discrete translation ($\Theta T_{1/2}$) symmetry is classified by a $\mathbb{Z}_{2}$-valued topological invariant and characterized by quantized magnetoelectric response. Here we demonstrate by explicit calculation in model $\mathbb{Z}_{2}$ topological insulator thin-films that whereas the magnetoelectric response is localized at the surface in the $\Theta$ symmetry (nonmagnetic) case, it is nonuniversally partitioned between surface and interior contributions in the $\Theta T_{1/2}$ (antiferromagnetic) case, while remaining quantized. Within our model the magnetic field induced polarization arises entirely from an anomalous ${\cal N}=0$ Landau level subspace within which the projected Hamiltonian is a generalized Su-Schrieffer-Heeger model whose topological properties are consistent with those of the starting 3D model. We identify a new connection between the ground-state geometry of that 3D model and surface-interior partitioning in thin films.
I Introduction
The realization that material properties can depend qualitatively [1, 2] on topological data encoded in the crystal-momentum dependence of its occupied electronic eigenspace projectors has been an important achievement in modern condensed matter physics. The physical import of this mathematical structure was first recognized in theoretical studies [3, 4] of the integer quantum Hall effect (IQHE) in which the Hall conductivity of two-dimensional band insulators was identified as the (quantized) Chern number of the vector bundle of occupied (electronic) Bloch states over the Brillouin zone (BZ) torus [5]. Subsequently other, more subtle, topological invariants of band insulating ground states have been found to be physically relevant. Among these, three-dimensional (3D) band insulators with time-reversal symmetry (TRS, denoted as $\Theta$) or generalized TRS ($\Theta T_{1/2}$), which have a quantized magnetoelectric response [6, 7, 8],
@@ I Introduction
have been a prominent focus. In Eq. (1) superscript indices identify Cartesian components, ${\mathbf{P}}$ and ${\mathbf{M}}$ are the (electronic) polarization and orbital magnetization, respectively, and ${\mathbf{E}}$ and ${\mathbf{B}}$ are uniform dc electric and magnetic fields, respectively. In both nonmagnetic ($\Theta$ symmetric) and antiferromagnetic (AFM) ($\Theta T_{1/2}$ symmetric [9]) cases, $\alpha=0$ or $e^{2}/2hc$ (mod $e^{2}/hc$) and is determined, in part, by the $\mathbb{Z}_{2}$ classification of the ground state [6, 7, 8]. Indeed, an explicit expression has been proposed [6] that relates $\alpha$ to the bulk BZ integral of the Chern-Simons 3-form in the space of occupied Bloch states. The value of $\alpha$ implied by this bulk expression ($\alpha_{\text{CS}}$), quoted explicitly below, is determined modulo $e^{2}/hc$ by the $\mathbb{Z}_{2}$ index of the vector bundle of occupied Bloch states [6, 9, 10, 11].
The relationship between $\alpha_{\text{CS}}$ and the physical magnetoelectric response ($\alpha_{\text{me}}$) of finite volume $\mathbb{Z}_{2}$ topological insulators (TIs) with surfaces is nontrivial [8, 12, 13, 14, 15]. In nonmagnetic thin films  [^1] We exclude systems with an integer quantum Hall effect, and distinguish surface and bulk of thin films. We therefore restrict our attention to films with even layer number greater than or equal to four. with surface magnetic dopants that have nonzero surface-normal magnetization components that are opposite on top and bottom surfaces, $\alpha_{\text{me}}$ approaches $\alpha_{\text{CS}}$ in the thick-film limit. Here we demonstrate by explicit calculation that whereas the quantized magnetoelectric response is localized at the surface in the nonmagnetic case, it is nonuniversally partitioned between surface and interior contributions in the AFM case. In both cases the response of the (electronic) charge density to magnetic field is local in the sense that it occurs only in the spatial vicinity of static magnetic moments that break $\Theta$ symmetry. The difference is that these are present throughout the film interior in the AFM case. The dipole moment of the charge density that is induced by the magnetic field in the interior region of an AFM film is nonuniversal, but for sufficiently thick films we find that the sum of interior and surface contributions to the dipole moment is quantized (mod $e^{2}/hc$) at either $\alpha_{\text{me}}=0$ or $e^{2}/2hc$ depending on the $\mathbb{Z}_{2}$ classification of the bulk ground state; in AFM (nonmagnetic) thin films (with magnetic surface dopants) $\Theta T_{1/2}$ ($\Theta$) symmetry is broken and therefore, e.g., $\alpha_{\text{me}}=0$ cannot be protected by that symmetry, instead it is protected by the bulk $\Theta T_{1/2}$ ($\Theta$)-symmetry-induced topological classification of the ground state. This property, that the sum of 3D interior and surface contributions is in accord with 3D bulk topological constraints even when the separate contributions are nonquantized and nonuniversal, is reminiscent of the frequently discussed nonuniversal Hall current partition between interior and edge [17, 18, 19, 20] in the IQHE.
Our explicit calculations employ a model that can describe both nonmagnetic and AFM layered materials, and for each case can support ground states characterized by a trivial or nontrivial $\mathbb{Z}_{2}$ index. The model provides a realistic description of A-type AFM TIs, similar to the paradigmatic AFM TI MnBi2Te4 [21, 22, 23, 24, 25, 26, 27, 28, 29, 30], but is simplified in a way that is convenient for explicit inclusion of magnetic fields that are parallel to the stacking axis [31]. We use a continuum version of this model to calculate electronic charge distributions across thin films and a lattice regularized version to address the topological magnetoelectric response of bulk crystals. As illustrated in Fig. 1(a), our thin film calculations show that the magnetic field induced dipole moment arises entirely from the model’s anomalous $\mathcal{N}=0$ Landau level (LL) subspace, the dynamics of which are governed by a generalized Su-Schrieffer-Heeger (SSH) Hamiltonian. Because the $\mathcal{N}\neq 0$ LL eigenfunctions are symmetric across the thin film, as illustrated in Figs. 1(b) and 1(c), they do not contribute to the magnetic field induced dipole moment. The absence of a dipole moment in the interior of a nonmagnetic film is illustrated schematically in Fig. 1(b), and its presence in an AFM film in Fig. 1(c).
@@
II Anomalous Landau levels
An effective low-energy Hamiltonian for the electronic states in the layered materials of interest can be constructed [32, 33, 31] using Dirac cones, one associated with top and bottom surface of each layer, that are coupled by hybridizing states both within and between layers:
@@ II Anomalous Landau levels
Here we have chosen $\hat{\bm{z}}$ as the stacking axis, ${\mathbf{k}}_{\perp}=(k_{x},k_{y})$, $i,j\in\{0,\ldots,2N-1\}$ where $i=2(l-1)$ $(i=2l-1)$ labels states that are associated with the bottom (top) surface of layer $l\in\{1,\ldots,N\}$, $\mu,\gamma\in\{\uparrow,\downarrow\}$ are spin labels, $\Delta_{ij}$ are the intra- and interlayer hybridization parameters, $v_{{}_{D}}$ is the velocity of an isolated Dirac cone, and $m_{i}$ are mass terms that account for exchange coupling to static local magnetic moments (which we take to be aligned with the stacking direction). We will consider only even layer number $N$ so that AFM thin films do not have uncompensated moments, and in our model $m_{i}$ is a mass term at surface $i$ that accounts for exchange coupling to static local magnetic moments. We set $m_{i}=\sum_{l}J_{il}M_{l}$ with dimensionless magnetization $M_{l}=0$ if layer $l$ is nonmagnetic and ±$\pm 1$ if layer $l$ is magnetic (for antiferromagnetic configuration we take $M_{l}=(-1)^{l}$); the Dirac-like states at the top or bottom surface of layer $l$ are coupled to the static magnetization of that layer (via $J_{\text{S}}$) and to that of the adjacent layer nearest that surface (via $J_{\text{D}}$), and hybridized only to adjacent surfaces with hopping $\Delta_{\text{S}}$ for opposite surfaces of the same layer and $\Delta_{\text{D}}$ for the closest surface of the adjacent layer (see Fig. 1 of Ref. [31]). In thin film models the Dirac-like states of the outermost surfaces do not have $J_{\text{D}}$ or $\Delta_{\text{D}}$ coupling.
We account for an external magnetic field ${\mathbf{B}}=-B\hat{{\mathbf{z}}}$ though the usual prescription  [^2] See, e.g., Chapter 2 of Winkler [65]. of mapping a ${\mathbf{k}}\cdot{\mathbf{p}}$ model to an envelope function approximated (real-space) Hamiltonian that is then minimally coupled to the electromagnetic field: $\hbar{\mathbf{k}}\rightarrow-i\hbar\bm{\nabla}-(eBx/c)\hat{{\mathbf{y}}}$ in the Landau gauge. Here $e>0$ is the elementary charge unit. As described in Appendix A, we apply this recipe to Eq. (II) and seek energy eigenfunctions of the form $\Psi_{E,q_{y}}(x,y)=e^{iq_{y}y}\Phi_{E,q_{y}}(x)/\sqrt{L_{y}}$, where $\Phi_{E,q_{y}}(x)$ has $2N\times 2$ components that correspond with layer and spin label degrees of freedom, respectively. This results in the individual Dirac Hamiltonian $H_{D}$ of each surface being recast in terms of the LL raising and lowering (differential) operators $a^{\dagger}\equiv(\tilde{x}-\partial_{\tilde{x}})/\sqrt{2}$ and $a\equiv(\tilde{x}+\partial_{\tilde{x}})/\sqrt{2}$,
$H_{D}=\hbar v_{{}_{D}}(\hat{z}\times{\mathbf{\sigma}})\cdot{\mathbf{k}}_{\perp}\to\hbar\omega_{c}\begin{pmatrix}0&a^{\dagger}\\ a&0\end{pmatrix},$ (3)
@@
Below we calculate the electric dipole moment of finite thickness films by endowing the LL eigenfunctions of the quasi-2D model Eq. (II) with $z$-dependence by taking $C^{{\cal N},r}_{(i,\mu)}\rightarrow C^{{\cal N},r}_{(i,\mu)}\delta(z-z_{i})$, where $z_{i}\hat{{\mathbf{z}}}$ is the location of the $i$th surface with the center of the film taken as the origin. The macroscopic electronic polarization (dipole moment per volume) is:
@@
where $\Omega=A(z_{2N-1}-z_{0})$ is the volume of the finite sample, $A$ is the film cross sectional area and we have noted that only negative energy states are occupied at neutrality. Here the $\Psi_{E}({\mathbf{r}})$ are real-space energy eigenfunctions that have support only within the sample. Taking account of the LL degeneracy $A/(2\pi l_{B}^{2})$, we obtain
@@
The factor in round brackets is the areal electronic charge density on surface $i$ and $d$ is the thickness of the unit cell for each layer. Since $m_{i}$ is odd under the mirror transformation $z\to-z$ (see Appendix A.1), it yields a spin-polarization that is odd in $z$. In the ${\cal N}=0$ anomalous LL subspace, which contains only spin-up states, the mass terms give rise to a charge polarization. Because the block projected Hamiltonian is odd under the combined mirror and spin-reversal operations, the charge density contributed by the ${\cal N}\neq 0$ blocks is even under $z\to-z$ and there is no contribution to the dipole moment; the magnetoelectric response of our model comes entirely from the ${\cal N}=0$ block. Since the Hamiltonian of the ${\cal N}=0$ sector is a generalized SSH model with alternating hopping strengths $\Delta_{\text{S}}$ and $\Delta_{\text{D}}$, the quantized magnetoelectric coefficient $\alpha_{\text{me}}$ in our model follows from familiar analogous polarization properties of the SSH model. In the following, we detail how these properties depend on model parameters.
III Bulk phase diagram and polarization properties
@@
We can obtain the topological phase diagram of the bulk 3D Hamiltonian related to Eq. (II) by tracking the band gap evolution when periodic boundary conditions are applied in $\hat{{\mathbf{z}}}$ direction  [^5] Technically, the model presented in Eq. (II) is a quasi-2D low-energy effective ${\mathbf{k}}\cdot\bm{p}$ Hamiltonian. To generate a genuine 3D model for the electronic states in bulk, we need to replace the ${\mathbf{k}}\cdot\bm{p}$ basis states that the quasi-2D Hamiltonian is expressed in terms of with hybrid Wannier functions that are spatially localized in the ${\mathbf{z}}$ direction and the set of which contains one spin-up and one spin-down WF localized near each surface. We then apply periodic boundary conditions, by which we mean we take the model to describe an infinite chain of coupled-Dirac cone models (i.e., it involves a sum over all lattice vectors) in the ${\mathbf{z}}$-direction.. Our findings are shown in Fig. 2(a). As explained below, these assignments have been confirmed by calculating the Zak phase of a corresponding one-dimensional (1D) model and also by explicitly evaluating the topological magnetoelectric coefficient of a 3D lattice regularization of Eq. (II). The model parameters are chosen such that $\delta\equiv J_{\text{D}}/J_{\text{S}}\rightarrow 1$ corresponds to the nonmagnetic case, for which a topological phase transition occurs when $|\Delta_{\text{D}}|=\Delta_{\text{S}}$. In this case there is no magnetic-field dependent charge transfer between layers in the bulk as illustrated by the gray curve in Fig. 2(b). When $\delta\neq 1$ the interior layers are magnetized and magnetic-field dependent charge transfer does occur as illustrated by the green curve in Fig. 2(b). Note that in the AFM case the topological phase boundary depends on $J_{\text{S}}$, that interlayer charge transfer occurs in both the trivial ($\mathbb{Z}_{2}$-even) and nontrivial ($\mathbb{Z}_{2}$-odd) phases with magnitude dependent on model parameters, and that the rate of charge transfer with field jumps at the phase transition points, as shown in Figs. 2(c) and 2(d).
@@
Figure 2: Topological phase diagram and bulk charge redistribution. (a) Phase diagram of bulk AFM topological insulators vs. $\Delta_{\text{D}}/\Delta_{\text{S}}$ and $\delta\equiv J_{\text{D}}/J_{\text{S}}$ at $\Delta_{\text{S}}/J_{\text{S}}=1$. (The phase boundaries for $J_{\text{S}}=0.8\Delta_{\text{S}}$ and $J_{\text{S}}=1.5\Delta_{\text{S}}$ are plotted as purple and yellow dashed lines.) (b) Schematic magnetic-field induced charge density redistribution for nonmagnetic (gray curve) and AFM (green curve) TIs; (c), (d) Wave-function amplitude variation (defined in panel (b)) vs. $\Delta_{\text{D}}$ for various values of $\delta$ when $J_{\text{S}}=1.5\Delta_{\text{S}}$ and $J_{\text{S}}=0.8\Delta_{\text{S}}$, respectively.
@@
The magnetic-field dependent charge transfer seems to suggest that the magnetoelectric responses of both $\mathbb{Z}_{2}$-even and $\mathbb{Z}_{2}$-odd phases are nonzero, but this is not the case. As we have described near Eq. (3) and detailed in Appendix A, the low-energy Hamiltonian in the 3D bulk can be block-diagonalized into subspaces that are uniquely identified by a LL index $\mathcal{N}$. As in finite samples, in this bulk model the magnetoelectric response comes entirely from the $\mathcal{N}=0$ LL subspace, within which the AFM Hamiltonian reduces to a family of identical 1D generalized SSH models – one for each state in the LL (see Appendix A.1):